Two impedances Z1 and Z2 are connected in parallel. The first branch takes a leading

current of 16 A and has a resistance of 5 Ω, while the second branch takes a lagging

current at 0.8pf. The applied voltage is 100+j200 V and the total power is 5 kW. Find

branch impedances, total circuit impedance, branch currents and total circuit current.


1
Expert's answer
2021-02-22T09:46:00-0500

Let, Current in the first branch be "I_1 = 16A"


Resistance in the first branch be "R_1 = 5\Omega"


Applied voltage be "V = 100+j200."


Total Power "P = 5 kW"


and, let the total current be I


We know,

"P=VI"


"5000 = (100+j200)I"


"I = \dfrac{5000}{100+j200}"


"I = \dfrac{(5000)(100-j200)}{50000}"


"I = 10-j20 A"


Impedance"Z_1" in the first branch can be calculated as "\dfrac{100+j200}{16}"


"Z_1" = "6.25+j12.5 \Omega"


Current "I_2" in the second branch can be calculated as,

"I_2 = I - I_1"


"I_2 = 10-j20-16"


"I_2 = -6-j20 A"


Impedance "Z_2" in second branch can be calculated as "\dfrac{100+j200}{-6-j20}"

"Z2 = -10.54+j1.83 \Omega"